# What Lurks Within? Concept Auditing for Shared Diffusion Models at Scale

### v1

@@ Abstract.
Diffusion models (DMs) have revolutionized text-to-image generation, enabling the creation of highly realistic and customized images from text prompts. With the rise of parameter-efficient fine-tuning (PEFT) techniques like LoRA, users can now customize powerful pre-trained models using minimal computational resources. However, the widespread sharing of fine-tuned DMs on open platforms raises growing ethical and legal concerns, as these models may inadvertently or deliberately generate sensitive or unauthorized content, such as copyrighted material, private individuals, or harmful content. Despite the increasing regulatory attention on generative AI, there are currently no practical tools for systematically auditing these models before deployment.
In this paper, we address the problem of concept auditing: determining whether a fine-tuned DM has learned to generate a specific target concept. Existing approaches typically rely on prompt-based input crafting and output-based image classification but suffer from critical limitations, including prompt uncertainty, concept drift, and poor scalability. To overcome these challenges, we introduce Prompt-Agnostic Image-Free Auditing (PAIA), a novel, model-centric concept auditing framework. By treating the DM as the object of inspection, PAIA enables direct analysis of internal model behavior, bypassing the need for optimized prompts or generated images. It integrates two key components: a prompt-agnostic strategy that mitigates prompt sensitivity by analyzing model behavior during late-stage denoising, and an image-free detection method based on conditional calibrated error, which compares the internal dynamics of a fine-tuned model against its base version. We evaluate PAIA on 320 controlled models trained with curated concept datasets and 690 real-world community models sourced from a public DM sharing platform, covering a wide range of concepts including celebrities, cartoon characters, videogame entities, and movie references. Evaluation results show that PAIA achieves over $90$% detection accuracy while reducing auditing time by 1840×18–40$\times$18 – 40 × compared to existing baselines. To our knowledge, PAIA is the first scalable and practical solution for pre-deployment concept auditing of diffusion models, providing a practical foundation for safer and more transparent diffusion model sharing.
Diffusion Models, Concept Auditing, Model Safety, Generative AI
copyright: acmlicensedjournalyear: 2025
@@ 1. Introduction
This shift has catalyzed a thriving ecosystem of community-driven model customization. Users can now fine-tune and distribute their models with minimal technical expertise, aided by user-friendly toolkits (automatic1111, ; comfy_website, ) and supported by sharing platforms such as CivitAI (civitai_website, ), HuggingFace (huggingface_website, ), and SeaArt (seaart_website, ). These platforms host tens of thousands of customized DMs, covering a broad spectrum of visual concepts, artistic styles, and application domains. While this democratization has opened new frontiers in creativity and accessibility, it has also introduced significant risks. Fine-tuned DMs can be misused or repurposed to generate inappropriate, offensive, or legally problematic content. Studies have documented instances where models replicate copyrighted characters (zhang2023copyright, ; shan2023glaze, ; zhu2024watermark, ), impersonate real individuals via deepfakes (hao2024doesn, ), or produce not safe for work (NSFW) content (zhang2024generate, ; Yang2023SneakyPromptJT, ; Chin2023Prompting4DebuggingRT, ).
Despite growing concerns, technical oversight remains minimal. Current auditing practices on public model hubs rely heavily on user-supplied tags with little automation or verification. For example, CivitAI (civitai_website, ), one of the most widely used platforms, relies primarily on user-provided metadata to flag models, such as “mature content” or depictions of “real people,” which are loosely defined, inconsistently applied, and easily circumvented. The absence of standardized review procedures or systematic validation has already led to legal consequences. The community reports and discussions suggest that CivitAI has received takedown requests related to the unauthorized use of copyrighted or personal content due to the models trained on proprietary content [^1] CivitAI Facilitates the Use of Stolen Intellectual Property, https://luddite.pro/civitai-facilitates-use-stolen-intellectual-property [^2] Square Enix may have filed a DMCA takedown notice with CivitAI, https://www.reddit.com/r/StableDiffusion/comments/11zubsj/it_appears_that_someone_acting_on_behalf_of.
@@
To mitigate such risks, some efforts have introduced proactive defenses, such as built-in safety filters designed to block harmful content by detecting unsafe prompts or screening generated images (Rando2022RedTeamingTS, ; Yang2023SneakyPromptER, ). However, their effectiveness is limited. These filters are often implemented as optional add-ons and can be easily disabled after model download, leaving the core DM unrestricted. More importantly, both these reactive and proactive defenses suffer from a shared limitation: they rely on observable behaviors, prompts and outputs, that are unstable, easy to manipulate, and difficult to verify at scale. These limitations call for a deeper shift in auditing methodology: rather than focusing on what a model produces, we ask a more direct and scalable question - what lurks within?
@@
In this work, we study the problem of concept auditing: given a specific concept or a curated set of concepts, such as copyrighted characters, company logos, or celebrity identities, can a shared, fine-tuned diffusion model generate content that embodies those concepts? Rather than attempting to flag all forms of inappropriate content, we focus on concept-specific auditing at scale. This reflects real-world enforcement needs, where moderation and legal action often target clearly defined entities tied to intellectual property. To keep the scope practical and aligned with real-world use cases, we focus on concepts that are visually distinctive and often subject to content moderation or IP protection. These include individual celebrity identities (e.g., “Taylor Swift”), characters from well-known cartoons (e.g., “Muppets”), and entities from games and movies. Our framework is designed to accommodate this diversity through an example-based approach: a concept is considered present if the model can generate recognizable outputs aligned with a small set of reference examples of the concept. This flexible definition allows us to support a broad range of concept granularities without relying on a rigid taxonomy. A detailed discussion of how we define and structure concept scopes for auditing can be found in Section 8.1.
@@
A common approach to this problem is to audit models through their outputs: by crafting prompts that might trigger the concept of interest, generating images from the model, and using a classifier to determine whether the concept appears, as shown in Figure 1. While, this output-driven pipeline has become the dominant strategy, it relies on two fragile assumptions: first, that effective prompts can be reliably discovered to generate the target concept; and second, that external detectors can accurately identify whether the generated output matches that concept. In practice, both assumptions often fail for community-tuned models. A key issue is trigger uncertainty: identifying prompts that consistently activate a target concept is inherently difficult due to the vast, discrete, and ambiguous nature of the prompt space. Optimization-based methods such as adversarial or reinforcement learning often converge on unnatural or semantically misaligned phrases, undermining their reliability. Even when triggers are found, the assumption of accurate detection is weakened by concept drift (widmer1996learning, ; gama2014survey, ): external classifiers or CLIP-based detectors are typically trained on natural images and often perform poorly on synthetic outputs, especially when prompt probing introduces distributional shifts that degrade both detection accuracy and optimization guidance (zhang2024generate, ; wen2024hard, ). These problems are further compounded by limited scalability: discovering effective prompts and verifying them via image generation and classification is computationally intensive, often requiring hundreds of iterations per model.
@@
These limitations highlight a fundamental bottleneck in current auditing pipelines: their reliance on observable behaviors — prompts and outputs — that are inherently unstable, costly to evaluate, and easily manipulated. To overcome these issues, we propose a fundamentally different approach grounded in a shift of perspective: rather than auditing DMs by adjusting their inputs or analyzing their outputs, we treat the model itself as the source of truth. This leads to a new direction in concept auditing, determining what a model has actually learned by examining its internal behavior, rather than what it happens to generate.
@@
While prompt-agnosticism reduces reliance on brittle input probing, current output-based detection methods remain limited by their dependence on image generation and external classifiers, leading to concept drift and are prohibitively expensive at scale. To address this, we introduce our second strategy: an image-free design that further reinforces our model-centric paradigm. Instead of evaluating model outputs, we directly assess the model’s internal denoising dynamics. Specifically, we propose a metric called conditional calibrated error, which quantifies the behavioral discrepancy between a fine-tuned model and its corresponding base model when processing concept-relevant inputs. Notably, the base models are typically accessible in real-world deployments, because LoRA fine-tuning is designed to be modular, which requires the base model to be loaded alongside the LoRA weights at inference time. This makes the base model a natural and reliable reference point for behavioral comparison. By comparing internal activations directly, we can isolate fine-tuning effects and detect concept learning, without image sampling or reliance on noisy supervision.
Together, these strategies offer two key advantages. First, they enhance robustness by grounding auditing in the model’s own training dynamics, avoiding false positives caused by prompt misalignment or miscalibrated detectors. Second, they improve efficiency and scalability by operating entirely within the model’s latent space—eliminating the need for prompt optimization, image generation, or downstream classification. Crucially, the prompt-agnostic and image-free designs are mutually reinforcing: by shifting the focus away from unstable input-output behavior and toward internal model representations, they enable a principled and scalable approach to concept auditing. We operationalize this model-centric perspective in a unified framework: Prompt-Agnostic Image-Free Auditing (PAIA), designed to efficiently and effectively determine whether a fine-tuned DM can generate a given target concept. To the best of our knowledge, PAIA represents the first practical, scalable, and systematically validated solution for concept auditing in fine-tuned DMs. Our major contributions are summarized below.
@@
  • Extensive Experimental Validation in the Wild. We conduct comprehensive experiments to evaluate PAIA on both controlled and real-world settings. In our controlled evaluation, we fine-tune $320$ DMs, each on a specific target concept, where a concept refers to a recognizable visual entity such as an individual celebrity or a cartoon character. This dataset includes $50$ celebrity identities and $10$ cartoon characters. For real-world evaluation, we collect $690$ community-shared models from the Civitai platform, spanning $174$ celebrities, $145$ cartoon characters, $192$ videogame-related entities, and $179$ movie-based concepts. Across both settings, PAIA consistently achieves high accuracy (over $90$%) and efficiency (1840×18-40$\times$18 - 40 × speedup), significantly outperforming existing baselines. To the best of our knowledge, this constitutes the first large-scale, systematic evaluation of concept auditing for fine-tuned DMs.
@@ 2.1. Diffusion Models (DMs)
DMs have emerged as one of the most effective models for image generation (croitoru2023diffusion, ; ruiz2023dreambooth, ; zhang2023adding, ), powering commercial image-generation applications, such as Stable Diffusion (rombach2021highresolution, ), DALL-E 3 (betker2023improving, ), and MidJourney (borji2022generated, ). Conceptually, the diffusion process can be described as a stochastic, iterative procedure in which noise is gradually introduced into an image until it becomes indistinguishable from pure noise. Specifically, in the diffusion process, given an image $\bm{x}_{0}$, a time step $0\leq t\leq T$, and a white noise vector $\bm{\epsilon}_{t}\sim\mathcal{N}(0,\mathbf{I})$, a noisy image at time step $t$, $\mathbf{x}_{t}$, is generated
$\mathbf{x}_{t}=\sqrt{\overline{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\overline{\alpha}_{t}}\bm{\epsilon}_{t},$
where $\overline{\alpha}_{t}$ represents a noise scheduling factor controlling the amount of noise injected at time step $t$. The noise schedule is designed to smoothly transform the image from the original $\mathbf{x}_{0}$ to an almost pure noise distribution, $\mathbf{x}_{T}$, as $t\to T$.
@@
To reduce computing for high-resolution generation, Latent Diffusion Models (LDMs) (rombach2021highresolution, ) operate in a lower-dimensional latent space $\mathbf{z}_{t}$ and apply the same training objective.
Text-to-Image (T2I) generation: T2I generation leverages textual input, i.e., prompts, to guide the image generation process (croitoru2023diffusion, ; ruiz2023dreambooth, ; zhang2023adding, ). This begins by encoding the textual prompt into a prompt embedding $\mathbf{p}$ via a text encoder. The prompt embedding $\mathbf{p}$ is then incorporated into a DM ${\bm{\epsilon}}_{W}(\mathbf{z}_{t},\mathbf{p})$, conditioning the image generation process on the textual description. This integration is facilitated through a cross-attention mechanism. Specifically, text features are encoded and projected into a key vector $K$ and a value vector $V$ using linear projection matrices $W_{K}$ and $W_{V}$, respectively. Simultaneously, the latent image features are projected into a query vector $Q$ using another projection matrix $W_{Q}$. A cross-attention map $S$ is computed as
@@
During inference, the T2I generation process involves both conditional denoising (where the prompt embedding is set to $\mathbf{p}$) and unconditional denoising (where the prompt embedding is encoded on a null character, i.e., $\mathbf{p}=\varnothing$). To balance the impact of $\mathbf{p}$ on a generated image, the predicted noise at time step $t$ is calculated as
(5) ${\bm{\epsilon}}_{W}(\mathbf{z}_{t},\mathbf{p})={\bm{\epsilon}}_{W}(\mathbf{z}_{t},\varnothing)+\eta\left({\bm{\epsilon}}_{W}(\mathbf{z}_{t},\mathbf{p})-{\bm{\epsilon}}_{W}(\mathbf{z}_{t},\varnothing)\right),$
where $\eta>1$ is the guidance scale that controls the strength of text conditioning. A higher $\eta$ increases the influence of the text prompt, making the generated image more closely aligned with the textual description, at the expense of diversity. After the iterative denoising process, the final latent image representation $\mathbf{z}_{0}$ is decoded into a high-resolution image $\mathbf{x}_{0}$ using a pre-trained decoder.

### v2

@@ Abstract.
Diffusion models (DMs) have revolutionized text-to-image generation, enabling the creation of highly realistic and customized images from text prompts. With the rise of parameter-efficient fine-tuning (PEFT) techniques like LoRA, users can now customize powerful pre-trained models using minimal computational resources. However, the widespread sharing of fine-tuned DMs on open platforms raises growing ethical and legal concerns, as these models may inadvertently or deliberately generate sensitive or unauthorized content, such as copyrighted material, private individuals, or harmful content. Despite increasing regulatory attention on generative AI, there are currently no practical tools for systematically auditing these models before deployment.
In this paper, we address the problem of concept auditing: determining whether a fine-tuned DM has learned to generate a specific target concept. Existing approaches typically rely on prompt-based input crafting and output-based image classification but they suffer from critical limitations, including prompt uncertainty, concept drift, and poor scalability. To overcome these challenges, we introduce Prompt-Agnostic Image-Free Auditing (PAIA), a novel, model-centric concept auditing framework. By treating the DM as the object of inspection, PAIA enables direct analysis of internal model behavior, bypassing the need for optimized prompts or generated images. It integrates two key components: a prompt-agnostic strategy that mitigates prompt sensitivity by analyzing model behavior during late-stage denoising, and an image-free detection method based on conditional calibrated error, which compares the internal dynamics of a fine-tuned model against its base version. Our auditing setting assumes internal access to DMs, but does not require access to proprietary fine-tuning data or user prompts, an assumption aligned with how hosted platforms audit uploaded models. We evaluate PAIA on $320$ controlled models trained with curated concept datasets and $771$ real-world community models sourced from a public DM sharing platform, covering a wide range of concepts including celebrities, cartoon characters, videogame entities, and movie references. Evaluation results show that PAIA achieves over 9090% detection accuracy while reducing auditing time by $18$ - $40\times$ compared to existing baselines, and remains robust under adaptive attacks. To our knowledge, PAIA is the first scalable and practical solution for pre-deployment concept auditing of diffusion models, providing a practical foundation for safer and more transparent diffusion model sharing [^1] This is an extended version of the paper accepted at CCS 2025..
Diffusion Models; Concept Auditing; Generative AI
journalyear: 2025copyright: ccconference: Proceedings of the 2025 ACM SIGSAC Conference on Computer and Communications Security; October 13–17, 2025; Taipei, Taiwanbooktitle: Proceedings of the 2025 ACM SIGSAC Conference on Computer and Communications Security (CCS ’25), October 13–17, 2025, Taipei, Taiwandoi: 10.1145/3719027.3765052isbn: 979-8-4007-1525-9/2025/10ccs: Computing methodologies Computer visionccs: Security and privacy Digital rights management
@@ 1. Introduction
This shift has catalyzed a thriving ecosystem of community-driven model customization. Users can now fine-tune and distribute their models with minimal technical expertise, aided by user-friendly toolkits (automatic1111, ; comfy_website, ) and supported by sharing platforms such as Civitai (civitai_website, ), HuggingFace (huggingface_website, ), and SeaArt (seaart_website, ). These platforms host tens of thousands of customized DMs, covering a broad spectrum of visual concepts, artistic styles, and application domains. While this democratization has opened new frontiers in creativity and accessibility, it has also introduced significant risks. Fine-tuned DMs may be misused to generate inappropriate or legally problematic content. Studies have documented instances where models replicate copyrighted characters (zhang2023copyright, ; shan2023glaze, ; zhu2024watermark, ), impersonate real individuals via deepfakes (hao2024doesn, ), or produce inappropriate content (zhang2024generate, ; Yang2023SneakyPromptJT, ; Chin2023Prompting4DebuggingRT, ).
Despite growing concerns, technical oversight remains minimal. Current auditing practices on public model hubs rely heavily on user-supplied tags with little automation or verification. For example, Civitai (civitai_website, ), one of the most widely used platforms, relies primarily on user-provided metadata to flag models, such as depictions of “real people,” or “mature content,” which are loosely defined, inconsistently applied, and easily circumvented. The absence of standardized review procedures or systematic validation has already led to legal consequences. The community reports and discussions suggest that Civitai has received takedown requests related to the unauthorized use of copyrighted or personal content due to the models trained on proprietary content [^2] Civitai Facilitates the Use of Stolen Intellectual Property, https://luddite.pro/civitai-facilitates-use-stolen-intellectual-property [^3] Square Enix may have filed a DMCA takedown notice with Civitai, https://www.reddit.com/r/StableDiffusion/comments/11zubsj/it_appears_that_someone_acting_on_behalf_of.
@@
To mitigate such risks, some efforts have introduced proactive defenses, such as built-in safety filters designed to block harmful content by detecting unsafe prompts or screening generated images (Rando2022RedTeamingTS, ; Yang2023SneakyPromptER, ). However, their effectiveness is limited. These filters are often implemented as optional add-ons and can be easily disabled after model download, leaving the core DM unrestricted. More importantly, these defenses suffer from a shared limitation: they rely on observable behaviors, prompts and outputs, that are unstable, easy to manipulate, and difficult to verify at scale. These limitations call for a deeper shift in auditing methodology: rather than focusing on what a model produces, we ask a more direct and scalable question: what lurks within?
@@
In this work, we study the problem of concept auditing: given a specific concept or a curated set of concepts, such as copyrighted characters, company logos, or celebrity identities, can a shared, fine-tuned diffusion model generate content that embodies those concepts? Rather than attempting to flag all forms of inappropriate content, we focus on concept-specific auditing at scale. This reflects real-world enforcement needs, where moderation and legal action often target clearly defined entities tied to intellectual property. To keep the scope practical and aligned with real-world use cases, we focus on concepts that are visually distinctive and often subject to content moderation or IP protection. These include individual celebrity identities (e.g., “Taylor Swift”), characters from well-known cartoons (e.g., “Muppets”), and entities from games and movies. Our framework is designed to accommodate this diversity through an example-based approach: a concept is considered present if the model can generate recognizable outputs aligned with a small set of reference examples of the concept. This flexible definition allows us to support a broad range of concept granularities without relying on a rigid taxonomy. Details of how we define and structure concept scopes for auditing can be found in Section 7.1.
@@
A common approach to this problem is to audit models through their outputs: by crafting prompts that might trigger the concept of interest, generating images from the model, and using a classifier to determine whether the concept appears, as shown in Figure 1. While this output-driven pipeline has become the dominant strategy, it relies on two fragile assumptions: first, that effective prompts can be reliably discovered to generate the target concept; and second, that external detectors can accurately identify whether the generated output matches that concept. In practice, both assumptions often fail for community-tuned models. A key issue is trigger uncertainty: identifying prompts that consistently activate a target concept is inherently difficult due to the vast, discrete, and ambiguous nature of the prompt space. Optimization-based methods such as adversarial or reinforcement learning often converge on unnatural or semantically misaligned phrases, undermining their reliability. Even when triggers are found, the assumption of accurate detection is weakened by concept drift (widmer1996learning, ; gama2014survey, ): external classifiers or CLIP-based detectors are typically trained on natural images and often perform poorly on synthetic outputs, especially when prompt probing introduces distributional shifts that degrade both detection accuracy and optimization guidance (zhang2024generate, ; wen2024hard, ). These problems are further compounded by limited scalability: discovering effective prompts and verifying them via image generation and classification is computationally intensive, often requiring hundreds of iterations.
@@
These limitations highlight a fundamental bottleneck in current auditing pipelines: their reliance on observable behaviors (i.e., prompts and their outputs) is inherently unstable, costly to evaluate, and easily manipulated. To overcome these issues, we propose a fundamentally different approach grounded in a shift of perspective: rather than auditing DMs by adjusting their inputs or analyzing their outputs, we treat the model itself as the source of truth. This motivates a new direction in concept auditing: determining what a model has actually learned by examining its internal behavior, rather than what it happens to generate.
@@
While prompt-agnosticism reduces reliance on brittle input probing, current output-based detection methods remain limited by their dependence on image generation and external classifiers, leading to concept drift, and are prohibitively expensive at scale. To address this, we introduce our second strategy: an image-free design that further reinforces our model-centric paradigm. Instead of evaluating model outputs, we directly assess the model’s internal denoising dynamics. Specifically, we propose a metric called conditional calibrated error, which quantifies the behavioral discrepancy between a fine-tuned model and its corresponding base model when processing concept-relevant inputs. Notably, the base models are typically accessible in real-world deployments, because LoRA fine-tuning is designed to be modular, which requires the base model to be loaded alongside the LoRA weights at inference time. This makes the base model a natural and reliable reference point for behavioral comparison. By comparing internal activations directly, we can isolate fine-tuning effects and detect concept learning, without image sampling or reliance on noisy supervision.
Together, the proposed strategies offer two key advantages. First, they enhance robustness by grounding auditing in the model’s own training dynamics, avoiding false positives caused by prompt misalignment or miscalibrated detectors. Second, they improve efficiency and scalability by operating entirely within the model’s latent space—eliminating the need for prompt optimization, image generation, or downstream classification. Crucially, the prompt-agnostic and image-free designs are mutually reinforcing: by shifting the focus away from unstable input-output behavior and toward internal model representations, they enable a principled and scalable approach to concept auditing. We operationalize this model-centric perspective in a unified framework: Prompt-Agnostic Image-Free Auditing (PAIA), designed to efficiently and effectively determine whether a fine-tuned DM can generate a given target concept. In this paper, we assume the auditor has internal access to the uploaded DMs (parameters), but without access to proprietary fine-tuning data or user prompts, which aligns with standard auditing workflows on hosted model hubs. To the best of our knowledge, PAIA represents the first practical, scalable, and systematically validated solution for concept auditing in fine-tuned DMs. Our major contributions are summarized below.
@@
  • Extensive Experimental Validation in the Wild. We conduct comprehensive experiments to evaluate PAIA on both controlled and real-world settings. In our controlled evaluation, we fine-tune $320$ DMs, each on a specific target concept, where a concept refers to a recognizable visual entity such as an individual celebrity or a cartoon character. This dataset includes $50$ celebrity identities and $10$ cartoon characters. For real-world evaluation, we collect $771$ community-shared models from the Civitai platform, spanning $174$ celebrities, $145$ cartoon characters, $192$ videogame-related entities, and $179$ movie-based concepts, along with $81$ rare concepts. Across both settings, PAIA consistently achieves high accuracy (over 9090%), efficiency ($18-40\times$ speedup), and remains robust under adaptive attacks, significantly outperforming existing baselines. To the best of our knowledge, this constitutes the first large-scale, systematic evaluation of concept auditing for fine-tuned DMs.
@@ 2.1. Diffusion Models (DMs)
DMs have emerged as one of the most effective models for image generation (croitoru2023diffusion, ; ruiz2023dreambooth, ; zhang2023adding, ), powering commercial image-generation applications, such as Stable Diffusion (rombach2021highresolution, ), DALL-E 3 (betker2023improving, ), and MidJourney (borji2022generated, ). Conceptually, the diffusion process can be described as a stochastic, iterative procedure in which noise is gradually introduced into an image until it becomes indistinguishable from pure noise. Specifically, in the diffusion process, given an image $\bm{x}_{0}$, a time step $0\leq t\leq T$, and a white noise vector $\bm{\epsilon}_{t}\sim\mathcal{N}(0,\mathbf{I})$, a noisy image at time step $t$, $\mathbf{x}_{t}$, is generated $\mathbf{x}_{t}=\sqrt{\overline{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\overline{\alpha}_{t}}\bm{\epsilon}_{t},$ where $\overline{\alpha}_{t}$ represents a noise scheduling factor controlling the amount of noise injected at time step $t$. The noise schedule is designed to smoothly transform the image from the original $\mathbf{x}_{0}$ to an almost pure noise distribution, $\mathbf{x}_{T}$, as $t\to T$.
@@
To reduce computational complexity for high-resolution generation, Latent Diffusion Models (LDMs) (rombach2021highresolution, ) operate in a lower-dimensional latent space $\mathbf{z}_{t}$ and apply the same training objective.
Text-to-Image (T2I) generation: T2I generation leverages textual input, i.e., prompts, to guide the image generation process (croitoru2023diffusion, ; ruiz2023dreambooth, ; zhang2023adding, ). This begins by encoding the textual prompt into a prompt embedding $\mathbf{p}$ via a text encoder. The prompt embedding $\mathbf{p}$ is then fed into a DM ${\bm{\epsilon}}_{W}(\mathbf{z}_{t},\mathbf{p})$, conditioning the image generation process on the textual description. This integration is facilitated through a cross-attention mechanism. Specifically, text features are encoded into a key vector $K$ and a value vector $V$ using linear projection matrices $W_{K}$ and $W_{V}$, respectively. Simultaneously, the latent image features are projected into a query vector $Q$ using another projection matrix $W_{Q}$. A cross-attention map $S$ is computed as
@@
During inference, the T2I generation process involves both conditional denoising (where the prompt embedding is set to $\mathbf{p}$) and unconditional denoising (where the prompt embedding is encoded on a null character, i.e., $\mathbf{p}=\varnothing$). To balance the impact of $\mathbf{p}$ on a generated image, the predicted noise at time step $t$ is calculated as ${\bm{\epsilon}}_{W}(\mathbf{z}_{t},\mathbf{p})={\bm{\epsilon}}_{W}(\mathbf{z}_{t},\varnothing)+\eta\left({\bm{\epsilon}}_{W}(\mathbf{z}_{t},\mathbf{p})-{\bm{\epsilon}}_{W}(\mathbf{z}_{t},\varnothing)\right),$ where $\eta>1$ is the guidance scale that controls the strength of text conditioning. A higher $\eta$ increases the influence of the text prompt, making the generated image more closely aligned with the textual description, at the expense of diversity. After the iterative denoising process, the final latent image representation $\mathbf{z}_{0}$ is decoded into a high-resolution image $\mathbf{x}_{0}$ using a pre-trained decoder.